# Extending Douglas–Rachford Splitting for Convex Optimization

### v1

@@ 4.1 Lyapunov Analysis
Unless stated otherwise, the following is valid for both $i=1$ and $i=2$. From Propositions˜4.1 and 4.1 we conclude after telescoping (LE) that, for all $K\in\mathbb{N}$:
$0\leq\sum_{k=0}^{K}\left(R_{k}^{(i)}+\theta\alpha I_{k}\right)\leq V_{0}^{(i)}-V_{K+1}^{(i)}\leq V_{0}^{(i)},$
where we also used $\alpha,\theta>0$ for all $(\alpha,\beta,\theta)\in S^{(i)}$. By letting $K\to\infty$, we conclude that (i) holds.
Moreover, since all coefficients in front of the terms in $R_{k}^{(i)}$ are nonnegative for $(\alpha,\beta,\theta)\in S^{(i)}$, since the Bregman-type terms are nonnegative by Lemma˜2.3, and since $\tau=\alpha/\beta$ and $\theta(2-\theta)$ are strictly positive, we get that $R_{k}^{(i)}\to 0$ as $k\to\infty$ implies that $x^{k}_{1}-x^{k}_{2}\to 0$ as $k\to\infty$. By Definition˜4.1, we have that
@@
$0\leq V_{k+1}^{(i)}\leq V_{k}^{(i)}-R_{k}^{(i)}\leq V_{k}^{(i)}$
for all $k\in\mathbb{N}$, we conclude that $\left(V_{k}^{(i)}\right)_{k\in\mathbb{N}}$ is a sequence of nonincreasing nonnegative numbers and so $\left(V_{k}^{(i)}\right)_{k\in\mathbb{N}}$ converges. Since all terms of $V_{k}^{(i)}$ are nonnegative, we have in particular if $i=1$ that $\tau(1-\tau)\left\lVert x_{1}^{k}-x^{\star}\right\rVert^{2}$ is, so $((1-\tau)(x_{1}^{k}-x^{\star}))_{k\in\mathbb{N}}$ is bounded since $\tau>0$. Using this, since also
@@
is bounded and combining with (ii), this shows boundedness of $(z^{k}-z^{\star})_{k\in\mathbb{N}}$. If instead $i=2$, by a similar argument, we have that
@@
is bounded, which combined with (ii) shows that $(z^{k}-z^{\star})_{k\in\mathbb{N}}$ is bounded.
By nonexpansiveness of resolvents, [BauschkeCombettes2017, Proposition 2.38], the update of Algorithm˜1, and the fixed-point set in Definition˜3.1, this implies that $(x_{1}^{k}-x^{\star})_{k\in\mathbb{N}}$ and $(x_{2}^{k}-x^{\star})_{k\in\mathbb{N}}$ are bounded. Therefore, $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ is bounded and by Definition˜4.1 we conclude that $(u_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ are bounded, so (iii) holds.
Using $u^{\star}\in\mathcal{H}$ as in Definition˜4.1, we conclude that
@@
as $k\to\infty$, where the first equality is due to Lemma˜2.2, the inequality is due to the monotonicity of $B$ since $u_{2}^{k}\in B(x_{2}^{k})$ and $u^{\star}=B(x^{\star})$, and the limit is a consequence of boundedness of $(x_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ combined with (ii). We conclude that $D_{f}(x^{\star},x_{1}^{k},u_{1}^{k})\to 0$ and $D_{f}(x_{1}^{k},x^{\star},u^{\star})\to 0$ as $k\to\infty$. Combining this with the fact that $\left(V_{k}^{(i)}\right)_{k\in\mathbb{N}}$ converges, we get in light of the definition of $V_{k}^{(i)}$ in Definition˜4.2 that (iv) holds. ∎
@@
We begin by showing that all weak sequential cluster points of $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ are elements of $\operatorname{Fix}(\partial f,B,\alpha,\beta)$. To that end, suppose that $(\bar{x}_{1},\bar{x}_{2},\bar{z})\in\mathcal{H}^{3}$ is a weak sequential cluster point of $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$. By Proposition˜4.2LABEL:*prp::analysis_ii, we get that $\bar{x}_{1}=\bar{x}_{2}$. Let $\bar{x}\coloneqq\bar{x}_{1}$. With the definition of $u_{2}^{k}$ in Definition˜4.1, we get that the subsequence $(x^{k_{n}}_{1},u^{k_{n}}_{2})\rightharpoonup\left(\bar{x},\frac{1}{\alpha}\left(\bar{x}-\bar{z}\right)\right)$ as $n\to\infty$. Moreover
@@
Let $i\in\{1,2\}$ be such that $(\alpha,\beta,\theta)\in S^{(i)}$. Let $Q_{V}^{(i)}\in\mathbb{S}^{3}$ be the matrix from Remark˜4.4, see also Appendix˜A for the explicit form. By Proposition˜4.2LABEL:*prp::analysis_iv, both $\left(\mathcal{Q}\left(Q_{V}^{(i)},v^{k}-\bar{v}\right)\right)_{k\in\mathbb{N}}$ and $\left(\mathcal{Q}\left(Q_{V}^{(i)},v^{k}-\hat{v}\right)\right)_{k\in\mathbb{N}}$ converge. Since
@@
we get that $\left(\left\langle v^{k},Q_{V}^{(i)}\left(\bar{v}-\hat{v}\right)\right\rangle\right)_{k\in\mathbb{N}}$ converges as well. Therefore, passing along the subsequences $(v^{k_{n}})_{n\in\mathbb{N}}$ and $(v^{\ell_{n}})_{n\in\mathbb{N}}$ gives that
@@
$\mathcal{Q}\left(Q_{V}^{(i)},\bar{v}-\hat{v}\right)=0.$ (8)
@@
where $1-\tau\geq 0$ for $(\alpha,\beta,\theta)\in S^{(1)}$. If $i=2$, then (8) equals
@@
where $4\tau-\theta$, $4-\theta$ and $\tau-1$ are nonnegative for $(\alpha,\beta,\theta)\in S^{(2)}$.
@@
Therefore, $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ is bounded and every weak sequential cluster point of $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ belongs to $\operatorname{Fix}(\partial f,B,\alpha,\beta)$, and so by [BauschkeCombettes2017, Lemma 2.47], the sequence $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ converges weakly to a point in $\operatorname{Fix}(\partial f,B,\alpha,\beta)$. ∎
@@ 4.2 Counterexamples
In the previous section, we have shown convergence of Algorithm˜1 over the parameter region $S^{(1)}\cup S^{(2)}$. With the following counterexamples we show that this region is tight, in the sense that Algorithm˜1 fails to be unconditionally convergent over $\partial\mathcal{F}\times\partial\mathcal{F}$ for parameters $(\alpha,\beta,\theta)$ outside of $S^{(1)}\cup S^{(2)}$.
@@
First, let $f=0\in\mathcal{F}$ and $g=\iota_{\{0\}}\in\mathcal{F}$. Note that $\operatorname*{zer}(\partial f+\partial g)=\{0\}$. Let $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ be generated by Algorithm˜1 applied to $(\partial f,\partial g,z^{0},\alpha,\beta,\theta)$ given some initial point $z^{0}\in\mathcal{H}\setminus\{0\}$. Note that such a $z^{0}$ exists since $\dim\mathcal{H}\geq 1$. Then,
@@
for every $k\in\mathbb{N}$. Therefore, we conclude in particular that $\theta\in(0,2)$.
Likewise, let us repeat the above argument applied for $f=\iota_{\{0\}}\in\mathcal{F}$ and $g=0\in\mathcal{F}$ with again $\operatorname*{zer}(\partial f+\partial g)=\{0\}$. Then
@@
Therefore, we conclude in particular that $\theta\in(0,2\alpha/\beta)$.
@@
Proposition˜4.4 shows that $S(\partial\mathcal{F}\times\partial\mathcal{F})\subset S^{(1)}\cup S^{(2)}$. Therefore,
@@ 5 Convergence Rates
We define the Lagrangian functions $\mathcal{L}:\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{-\infty,+\infty\}$ and $\bar{\mathcal{L}}:\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{-\infty,+\infty\}$ as
$\mathcal{L}(x,u)$ $\coloneqq f(x)+\langle x,u\rangle-g^{*}(u)$
$\bar{\mathcal{L}}(x,u)$ $\coloneqq g(x)+\langle x,u\rangle-f^{*}(u)$
and the primal-dual gap functions $\mathcal{D}_{x^{\star},-u^{\star}}:\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{+\infty\}$ and $\bar{\mathcal{D}}_{x^{\star},u^{\star}}:\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{+\infty\}$ as
@@
for all $(x,u)\in\mathcal{H}^{2}$, we have by convexity of $f$ and $g$ and by ˜9 combined with Lemma˜2.3 that $\mathcal{D}_{x^{\star},-u^{\star}}$ and $\bar{\mathcal{D}}_{x^{\star},u^{\star}}$ are convex and nonnegative.
@@
Let $(f,g)\in\mathcal{F}^{2}$ such that $\operatorname*{zer}(\partial f+\partial g)\neq\emptyset$ and let $(x^{\star},u^{\star})\in\mathcal{H}^{2}$ satisfy ˜9. Let $(\alpha,\beta,\theta)\in S^{(i)}$ for $i\in\{1,2\}$. For $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ generated by Algorithm˜1 applied to $(\partial f,\partial g,\alpha,\beta,\theta)$ and $(u_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ as in Definition˜4.1 we define the ergodic iterates
@@
For the ergodic iterates in Definition˜5.1, the nonnegative primal-dual gap sequences
@@
We will use the following identity. If $h:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ is proper, closed and convex, $x,y\in\mathcal{H}$, $u\in\partial h(x)$, and $v\in\partial h(y)$, then
@@
$\quad-\left(h(x)+h^{*}(u)-\langle x,u\rangle\right)-\langle u,y\rangle+\langle x,v-u\rangle$
@@
since the expressions in the parentheses are zero by the Fenchel-Young equality.
By Proposition˜3.1, there exists some $z^{\star}\in\mathcal{H}$ such that $(x^{\star},x^{\star},z^{\star})\in\operatorname{Fix}(\partial f,\partial g,\alpha,\beta)$. Combining ˜10 with ˜11, and then using ˜12 for $f$ and $g$ we get that
@@
$\leq\frac{1}{K+1}\sum_{k=0}^{K}I_{k}\leq\frac{1}{\alpha\theta(K+1)}\sum_{k=0}^{K}(\alpha\theta I_{k}+R_{k}^{(i)})\leq\frac{V_{0}^{(i)}}{\alpha\theta(K+1)}$
for all $K\in\mathbb{N}$ and for $V_{k}^{(i)}$ and $R_{k}^{(i)}$ defined in Definition˜4.2 and $I_{k}$ defined in Definition˜4.3, where the first equality follows from ˜13, the second equality from Definition˜4.3, the second inequality from monotonicity of $\partial f$ and $\partial g$ since $(x_{1}^{k+1},u_{1}^{k+1}),(x^{\star},u^{\star})\in\operatorname*{gra}\partial f$ and $(x_{2}^{k+1},u_{2}^{k+1}),(x^{\star},-u^{\star})\in\operatorname*{gra}\partial g$ by Definition˜4.1, the third inequality from nonnegativity of $R_{k}^{(i)}$ (Proposition˜4.2), and the fourth inequality from telescope summation of the Lyapunov equality (7) in Proposition˜4.1 and nonnegativity of $V_{k}^{(i)}$ (Proposition˜4.2). By nonnegativity of the primal-dual gap functions this shows ergodic convergence with rate $\mathcal{O}(1/K)$. ∎
6 Extending ADMM, Chambolle–Pock, and Parallel Splitting
In this section, we apply Algorithm˜1 to different problem formulations and, by invoking Theorem˜3.3, arrive at new convergent variants of ADMM, the Chambolle–Pock method, and Douglas–Rachford splitting applied to a consensus problem formulation, sometimes referred to as parallel splitting.

### v2

@@ 4.1 Lyapunov Analysis
From Propositions˜4.1 and 4.2 we conclude after telescoping (LE) that, for all $K\in\mathbb{N}$:
$0\leq\sum_{k=0}^{K}\left(R_{k}+\theta\alpha I_{k}\right)\leq V_{0}-V_{K+1}\leq V_{0},$
where we also use that $\alpha\theta>0$ for all $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$. By letting $K\to\infty$, we conclude that (i) holds.
Let $i\in\{1,2\}$ such that $(\alpha,\beta,\theta)\in S^{(i)}$. By Proposition˜4.1 and (i), we have that $R_{k}\geq 0$ and $R_{k}\to 0$ as $k\to\infty$, respectively. If $i=1$, by using the expression for $R_{k}$ from Definition˜4.2, nonnegativity of $(1-\tau)$ and $2\alpha(1-\tau)$, nonnegativity of the Bregman-type term by Lemma˜2.3, and strict positivity of $\theta(2\tau-\theta)$, we get that $x^{k}_{1}-x^{k}_{2}\to 0$ as $k\to\infty$.
If $i=2$, an analogous argument holds. Indeed, by using the expression for $R_{k}$ from Lemma˜4.1, by nonnegativity of $(\tau-1)$ and $2\alpha(\tau-1)$, nonnegativity of the Bregman-type term by Lemma˜2.3, and strict positivity of $\tau\theta(2-\theta)$, we get that $x^{k}_{1}-x^{k}_{2}\to 0$ as $k\to\infty$.
By Definition˜4.1 we have that
@@
$0\leq V_{k+1}\leq V_{k}-R_{k}\leq V_{k}$
for all $k\in\mathbb{N}$, we conclude that $\left(V_{k}\right)_{k\in\mathbb{N}}$ is a sequence of nonincreasing nonnegative numbers and so $\left(V_{k}\right)_{k\in\mathbb{N}}$ converges.
If $(\alpha,\beta,\theta)\in S^{(1)}$, then since all terms of $V_{k}$, given by Definition˜4.2, are nonnegative, we have in particular that $\tau(1-\tau)\left\lVert x_{1}^{k}-x^{\star}\right\rVert^{2}$ is bounded, implying that $\left((1-\tau)(x_{1}^{k}-x^{\star})\right)_{k\in\mathbb{N}}$ is bounded since $\tau>0$. Since the sequence
@@
is bounded, this fact, combined with (ii), shows that $(z^{k}-z^{\star})_{k\in\mathbb{N}}$ is bounded.
If instead $(\alpha,\beta,\theta)\in S^{(2)}$, by a similar argument and using Lemma˜4.1, we have that
@@
is bounded, this fact, combined with (ii), shows that $(z^{k}-z^{\star})_{k\in\mathbb{N}}$ is bounded.
By nonexpansiveness of resolvents, see [BauschkeCombettes2017, Proposition 2.38], the definitions of $(x_{1}^{k})_{k\in\mathbb{N}}$ and of the fixed-point set in Definition˜3.1, and boundedness of $(z^{k}-z^{\star})_{k\in\mathbb{N}}$, we get that $(x_{1}^{k}-x^{\star})_{k\in\mathbb{N}}$ is bounded. Boundedness of $(x_{2}^{k}-x^{\star})_{k\in\mathbb{N}}$ then follows from Proposition˜4.2LABEL:*prp::analysis_ii. Therefore, $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ is bounded and by Definition˜4.1 we conclude that $(u_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ are bounded, so (iii) holds.
Using $u^{\star}\in\mathcal{H}$ as in Definition˜4.1, we get that
@@
as $k\to\infty$, where the first equality is due to Lemma˜2.2, the second inequality is due to the monotonicity of $B$ since $u_{2}^{k}\in B(x_{2}^{k})$ and $u^{\star}\in B(x^{\star})$, and the limit is a consequence of boundedness of $(x_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ combined with (ii). We conclude that $D_{f}(x^{\star},x_{1}^{k},u_{1}^{k})\to 0$ and $D_{f}(x_{1}^{k},x^{\star},u^{\star})\to 0$ as $k\to\infty$. Combining this with the fact that $\left(V_{k}\right)_{k\in\mathbb{N}}$ converges, we get in light of the expressions for $V_{k}$, as given in Definition˜4.2 and Lemma˜4.1, that (iv) holds. ∎
@@
We begin by showing that all weak sequential cluster points of $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ are elements of $\operatorname{Fix}(\partial f,B,\alpha,\beta)$. To that end, suppose that $(\bar{x}_{1},\bar{x}_{2},\bar{z})\in\mathcal{H}^{3}$ is a weak sequential cluster point of $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$, say $(x^{k_{n}}_{1},x^{k_{n}}_{2},z^{k_{n}})\rightharpoonup(\bar{x}_{1},\bar{x}_{2},\bar{z})$ as $n\to\infty$. By Proposition˜4.2LABEL:*prp::analysis_ii, we get that $\bar{x}_{1}=\bar{x}_{2}$. Let $\bar{x}\coloneqq\bar{x}_{1}$. With the definition of $u_{2}^{k}$ in Definition˜4.1, we get that the subsequence $(x^{k_{n}}_{1},u^{k_{n}}_{2})\rightharpoonup\left(\bar{x},\frac{1}{\alpha}\left(\bar{x}-\bar{z}\right)\right)$ as $n\to\infty$. Moreover
@@
Let $i\in\{1,2\}$ be such that $(\alpha,\beta,\theta)\in S^{(i)}$ and let $\bar{Q}_{V}^{(i)}\in\mathbb{S}^{3}$ be the matrix from Remark˜4.2. By Proposition˜4.2LABEL:*prp::analysis_iv, both $\left(\mathcal{Q}\left(\bar{Q}_{V}^{(i)},v^{k}-\bar{v}\right)\right)_{k\in\mathbb{N}}$ and $\left(\mathcal{Q}\left(\bar{Q}_{V}^{(i)},v^{k}-\hat{v}\right)\right)_{k\in\mathbb{N}}$ converge. Since
@@
we get that $\left(\left\langle v^{k},\bar{Q}_{V}^{(i)}\left(\bar{v}-\hat{v}\right)\right\rangle\right)_{k\in\mathbb{N}}$ converges as well. Therefore, passing along the subsequences $(v^{k_{n}})_{n\in\mathbb{N}}$ and $(v^{\ell_{n}})_{n\in\mathbb{N}}$ gives that
@@
$\mathcal{Q}\left(\bar{Q}_{V}^{(i)},\bar{v}-\hat{v}\right)=0.$ (9)
@@
where $(1-\tau)$ is nonnegative for $(\alpha,\beta,\theta)\in S^{(1)}$. If $i=2$, then (9) equals
@@
where $(4\tau-\theta)$, $(4-\theta)$ and $(\tau-1)$ are nonnegative for $(\alpha,\beta,\theta)\in S^{(2)}$.
@@
Therefore, $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ is bounded and every weak sequential cluster point of $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ belongs to $\operatorname{Fix}(\partial f,B,\alpha,\beta)$, and so by [BauschkeCombettes2017, Lemma 2.46], the sequence $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ converges weakly, and since all weak sequential cluster points lie in $\operatorname{Fix}(\partial f,B,\alpha,\beta)$, we have that $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ converges weakly to a point in $\operatorname{Fix}(\partial f,B,\alpha,\beta)$. ∎
@@ 4.2 Counterexamples
In the previous section, we have shown weak convergence of Algorithm˜1 over the parameter region $S^{(1)}\cup S^{(2)}$. With the following counterexamples we show that this region is tight, in the sense that Algorithm˜1 fails to be unconditionally convergent over $\partial\mathcal{F}\times\partial\mathcal{F}$ for parameters $(\alpha,\beta,\theta)$ outside of $S^{(1)}\cup S^{(2)}$.
@@
First, let $f=0\in\mathcal{F}$ and $g=\iota_{\{0\}}\in\mathcal{F}$. Note that $\operatorname*{zer}(\partial f+\partial g)=\{0\}$. Let $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ be generated by Algorithm˜1 applied to $(\partial f,\partial g,z^{0},\alpha,\beta,\theta)$ given some initial point $z^{0}\in\mathcal{H}\setminus\{0\}$. Note that such an element $z^{0}$ exists since $\dim\mathcal{H}\geq 1$. Then
@@
for every $k\in\mathbb{N}$. We conclude in particular that $\theta\in(0,2)$.
Likewise, let us repeat the above argument for $f=\iota_{\{0\}}\in\mathcal{F}$ and $g=0\in\mathcal{F}$ where again $\operatorname*{zer}(\partial f+\partial g)=\{0\}$. Then
@@
for every $k\in\mathbb{N}$. We conclude in particular that $\theta\in(0,2\alpha/\beta)$.
@@
Proposition˜4.4 gives that $S(\partial\mathcal{F}\times\partial\mathcal{F})\subset S^{(1)}\cup S^{(2)}$. Therefore,
@@ 5 Convergence Rates
We define the Lagrangian functions $\mathcal{L}\colon\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{-\infty,+\infty\}$ and $\bar{\mathcal{L}}\colon\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{-\infty,+\infty\}$ as
$\mathcal{L}(x,u)$ $\coloneqq f(x)+\langle x,u\rangle-g^{*}(u),$
$\bar{\mathcal{L}}(x,u)$ $\coloneqq g(x)+\langle x,u\rangle-f^{*}(u).$
Whenever the expression $+\infty-(+\infty)$ arises, it is understood to take the value -$\infty$. This convention is adopted for definiteness, although this indeterminate case will not occur in what follows. We define the primal-dual gap functions $\mathcal{D}_{x^{\star},-u^{\star}}:\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{+\infty\}$ and $\bar{\mathcal{D}}_{x^{\star},u^{\star}}:\mathcal{H}^{2}\rightarrow\mathbb{R}\cup\{+\infty\}$ as
@@
for all $(x,u)\in\mathcal{H}^{2}$. From the convexity of $f$ and $g$, ˜10, and Lemma˜2.3, it follows that $\mathcal{D}_{x^{\star},-u^{\star}}$ and $\bar{\mathcal{D}}_{x^{\star},u^{\star}}$ are convex and nonnegative.
@@
Let $f,g$ be proper, closed, and convex functions, let $z^{0}\in\mathcal{H}$, and let $(\alpha,\beta,\theta)\in S^{(1)}\cup S^{(2)}$. For $\bigl((x^{k}_{1},x^{k}_{2},z^{k})\bigr)_{k\in\mathbb{N}}$ generated by Algorithm˜1 applied to $(\partial f,\partial g,z^{0},\alpha,\beta,\theta)$ with $(u_{1}^{k})_{k\in\mathbb{N}}$ and $(u_{2}^{k})_{k\in\mathbb{N}}$ as in Definition˜4.1, we define the ergodic iterates
@@
Following Definition˜5.1, let $(x^{\star},u^{\star})\in\mathcal{H}^{2}$ satisfy ˜10. Then the primal-dual gap sequences
@@
We will use the following identity. If $h:\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$ is proper, closed, and convex, $(x,y)\in\mathcal{H}^{2}$, $u\in\partial h(x)$, and $v\in\partial h(y)$, then
@@
$\quad-\left(h(x)+h^{*}(u)-\langle x,u\rangle\right)-\langle u,y\rangle-\langle x,v-u\rangle$
@@
since the expressions in the parentheses are zero by the Fenchel–Young equality.
By Proposition˜3.1, there exists some $z^{\star}\in\mathcal{H}$ such that $(x^{\star},x^{\star},z^{\star})\in\operatorname{Fix}(\partial f,\partial g,\alpha,\beta)$. Combining ˜11 with ˜12, and then using ˜13 for $f$ and $g$, we get that
@@
$\leq\frac{1}{K+1}\sum_{k=0}^{K}I_{k}\leq\frac{1}{\alpha\theta(K+1)}\sum_{k=0}^{K}(\alpha\theta I_{k}+R_{k})\leq\frac{V_{0}}{\alpha\theta(K+1)}$
for all $K\in\mathbb{N}$ and for $V_{k}$ and $R_{k}$ defined in Definition˜4.2 and $I_{k}$ defined in Definition˜4.3, where the first equality follows from ˜14, the second equality from Definition˜4.3, the second inequality from monotonicity of $\partial f$ and $\partial g$ since $(x_{1}^{k+1},u_{1}^{k+1}),(x^{\star},u^{\star})\in\operatorname*{gra}\partial f$ and $(x_{2}^{k+1},u_{2}^{k+1}),(x^{\star},-u^{\star})\in\operatorname*{gra}\partial g$ by Remark˜4.1, the third inequality from nonnegativity of $R_{k}$ (Proposition˜4.2), and the fourth inequality by summation and telescoping of the Lyapunov equality (LE) in Proposition˜4.1 and by nonnegativity of $V_{k}$ (Proposition˜4.2). By nonnegativity of the primal-dual gap functions, this shows ergodic convergence with rate $\mathcal{O}(1/K)$. ∎
6 Extending ADMM, Chambolle–Pock, and more
In this section, we apply Algorithm˜1 to different problem formulations and, by invoking Theorem˜3.3, arrive at new convergent variants of ADMM, the Chambolle–Pock method, and Douglas–Rachford splitting applied to a consensus problem formulation, sometimes referred to as parallel splitting. This section concludes with the derivation of a new version of the Generalized Alternating Projections method, allowing for overreflected projections.